# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v3

@@ V.2 UniPero Model
Wu et al. Wu et al. (2023a) utilized “modular development of deep potential” (ModDP) Wu et al. (2023b) to create a universal interatomic potential for perovskite oxides (UniPero) using a deep neural network with a self-attention mechanism. This potential spans 26 types perovskite oxides involving 200 components and 14 metal elements. All DFT calculations were performed with ABACUS. As depicted in Fig. 29(a), they first used DP-GEN to obtain a converged DPA-1 model Zhang et al. (2024c) for three-element systems such as PbTiO3, SrTiO3. Then, the converged dataset served as a starting point for DP-GEN to improve DPA-1 for four-element perovskite systems such as PbxSr1-xTiO3 and PbZrxTi1-xO3. Ultimately, the final DPA-1 model (UniPero) can describe six-element perovskite systems, including the ternary solid solution Pb(In1/2Nb1/2)O3-Pb(Mg1/3Nb2/3)O3–PbTiO3 (PIN-PMN-PT). UniPero functions as a universal interatomic potential, effectively modeling a wide range of perovskites through molecular dynamics simulations.
@@ V.3 DPA-Semi Model
Liu et al. Liu et al. (2024b) generated over 200,000 DFT data for 19 bulk semiconductors ranging from group IIB to VIA, namely, Si, Ge, SiC, BAs, BN, AlN, AlP, AlAs, InP, InAs, InSb, GaN, GaP, GaAs, CdTe, InTe, CdSe, ZnS, CdS. All of the DFT calculations were performed with the ABACUS v3.2 package with the Perdew-Burke-Ernzerhof (PBE) Perdew, Burke, and Ernzerhof (1996) exchange-correlation functional. The triple-zeta plus double polarization (TZDP) NAO basis sets were used for all DFT calculations. The atomic datasets are adopted as training data to generate an attention-based deep potential model using the DPA-1 method Zhang et al. (2024c), namely the DPA-Semi model. Fig. 30 shows the procedures for developing the DPA-Semi model.
@@ V.4 DPA-1, DPA-2 and DPA-3
The DPA-1 model Zhang et al. (2024c) is a universal machine learning interatomic potential that leverages the attention mechanism for molecular and materials modeling. Trained on the OC20 benchmark dataset and specialized materials systems, including aluminum-magnesium-copper alloys, high-entropy alloys, and solid-state electrolytes, this architecture achieves robust transferability across diverse atomic environments. With coverage spanning 56 elements in the periodic table, DPA-1 employs a pretraining-fine-tuning paradigm. The attention-based design supports first-principles accuracy while maintaining computational efficiency comparable to classical force fields.
@@
The DPA-2 model Zhang et al. (2024d) employs a multi-task learning framework that effectively integrates datasets from diverse DFT sources. This architecture substantially reduces the model’s dependence on data uniformity, enabling utilization of datasets generated by different methods. Notably, the pretraining phase of DPA-2 incorporates systematic learning of chemical compositions and spatial configurations, which dramatically reduces the demand for task-specific fine-tuning data in downstream applications.
@@
The recently developed DPA-3 Zhang et al. (2025) constructs a hierarchical graph neural network architecture based on line graph transformations, demonstrating substantial improvements in scaling law adherence and cross-dataset generalization capabilities. During the development of both DPA-2 and DPA-3 iterations, the ABACUS software served as a critical computational infrastructure, facilitating DFT calculations across diverse material datasets. Here are two examples.
@@
First, based on DFT calculations performed using ABACUS, we are developing general machine-learning-based potential models for alloys, covering 53 elements in the periodic table. The elements are Li, Na, K, Be, Mg, Ca, Sr, Sc, Y, La, Ti, Zr, Hf, V, Nb, Ta, Cr, Mo, W, Cu, Ag, Au, Zn, Cd, Mn, Re, Fe, Co, Ni, Os, Ir, Pt, Rh, Ru, Pd, Ce, Pr, Nd, Sm, Gd, Tb, Dy, Ho, Er, Tm, Lu, Al, Ga, In, Si, Ge, Sn, Pb, where bold ones are commonly used in structural alloys. The specific pseudopotentials utilized for these elements are listed in Ref. lin, 2024 and the calculations were performed efficiently by the ABACUS package on the Deep Computing Unit (DCU) hardware. Specifically, the DP-GEN and APEX workflows generated various structures Zhang et al. (2020c); Li et al. (2025), including perfect and perturbed crystal structures, vacancies, interstitials, and surfaces of metals and alloys. Some structures were selected through the concurrent learning workflow in DP-GEN and then labeled by ABACUS, resulting in a training dataset of $\sim$24,000 entries. By integrating previous training datasets from AI Square and the OpenLAM project ope (2024a), the latest general LAM for alloys achieved a root mean square error (RMSE) of $\sim$26 meV/atom for energies and $\sim$0.20 eV/Å for atomic forces across the 53 metals and their alloys. This general LAM for alloys demonstrates superior performance in predicting various properties, including lattice parameters, elastic constants, point defects, and surface formation energies, compared to other LAMs Batatia et al. (2024). ABACUS has proven to be a stable and reliable DFT workhorse for diverse structures covering these 53 elements.
@@
Second, during the construction of the dataset for the DPA-1 and DPA-2 models of high-pressure SuperHydrides, all labels were calculated using ABACUS with plane-wave basis sets. The thermodynamic stability of the discovered structures was confirmed by their energies above the convex hull, which was determined through structure optimization results from ABACUS. Dynamic stability was assessed using the phonon dispersion spectrum calculated by phonopy Togo et al. (2023); Togo (2023), with all forces calculated by ABACUS. Our DFT calculations were performed on DCU hardware, significantly enhancing the efficiency of SCF calculations compared to traditional CPUs. We obtained a dataset comprising 218,349 data frames, covering 29 elements and a pressure range of 150-250 GPa. Utilizing this data, the DPA-1 model achieved a training RMSE of 48.1 meV/atom for energy and 334.8 meV/Å for force. The DPA-2 model achieved a training RMSE of 55.2 meV/atom for energy and 298.8 meV/Å for force. On the testing dataset, the DPA-1 model yielded an RMSE of 37.6 meV/atom for energy and 171.4 meV/Å for force, while the DPA-2 model yielded an RMSE of 37.4 meV/atom for energy and 122.1 meV/Å for force. Given the large dataset required for the Superhydrides model, ABACUS proved to be the most economical DFT software available. The high-quality training results further attest to the stable performance of ABACUS.
@@ VI.1 DeePKS-kit
DeePKS-kit Chen et al. (2021, 2022a) is an innovative computational framework designed to bridge machine learning with quantum-mechanical simulations. The software integrates deep learning techniques with the Kohn-Sham equations to refine exchange-correlation functionals and can be used with ABACUS or PySCF Sun et al. (2020). By training neural networks on high-fidelity quantum chemistry data, DeePKS-kit dynamically corrects errors in conventional DFT approximations, enabling predictions of electronic properties with near-chemical-accuracy at a fraction of the computational cost.
By adopting both DeePKS-kit and ABACUS packages, we optimize the DeePKS model in two iterative steps. As shown in Fig. 32, $L(\omega)$ is the loss function. Loss function $L(\omega)$ can be composed of various energetic terms including the energy $E$, the atomic force $\mathbf{F}$, the stress tensor $\mathbf{\sigma}$ and the band gap $\varepsilon_{g}$ of the interested system The first step is the SCF step, where we fix parameters $\{\omega^{\ast}\}$ in the DeePKS model and solve the Kohn-Sham equation to obtain the ground state electronic wave function. The second step is the TRAIN step, where the electronic wave functions are fixed and the model parameters $\{\omega\}$are optimized. These two steps repeat until convergence is reached. Chen et al. (2021)
@@ VI.2 DeePMD-kit
DeePMD Zhang et al. (2018a, b) is a widely-used Wen et al. (2022) machine-learning molecular dynamics method based on the neural-network potential. ABACUS has an interface with the DeePMD-kit package Wang et al. (2018); Zeng et al. (2023), allowing its deployment as a molecular dynamics engine in executing DeePMD simulations. Trained on system energy, atomic forces, and lattice stress derived from DFT calculations, this neural network potential enables DeePMD to replicate the potential energy surface with first-principles accuracy. Since the neural network potential is computationally more efficient than the DFT method, DeePMD can simulate large systems at a long time scale, which is barely accessible for AIMD simulations.
@@ VI.3 DP-GEN
The DP-GEN framework has proven particularly effective for complex systems like multicomponent alloys Wu et al. (2021); Shi et al. (2024), chemical reaction pathways Zeng et al. (2021); Yang et al. (2022), and phase transitions under extreme conditions Gartner et al. (2020); Fu et al. (2023), where traditional sampling methods struggle to capture rare events or subtle energy landscapes. Open-source and modular by design, DP-GEN supports customization for specific materials or molecules, offering a versatile tool for accelerating the development of robust, quantum-mechanically informed potentials.
@@ VI.4 DeepH
The deep-learning density functional theory Hamiltonian (DeepH) method is a neural network approach based on equivariant graph neural networks for modeling the DFT Hamiltonian as a function of material structure Li et al. (2022b). Leveraging the sparsity of the Hamiltonian matrix under NAOs and its compatibility with Walter Kohn’s “quantum nearsightness principle" Kohn (1996), the DeepH method can learn from training data of small structures to infer electronic Hamiltonians of large structures, achieving high prediction accuracy with linear-scaling computational cost. The predicted Hamiltonian may be subsequently utilized for post-processing to evaluate properties including band structures, optical properties, and response properties from density functional perturbation theory, etc Li et al. (2022b); Gong et al. (2023); Li et al. (2024).
@@
The interface between ABACUS and DeepH was developed in 2022, shortly after the invention of the DeepH approach in 2021. The interface ensures compatibility between ABACUS and various versions of DeepH, including DeepH-pack Li et al. (2022b); DeepH , DeepH-E3 Gong et al. (2023); DeepH-E (3), and DeepH-2 Wang et al. (2024a). Among these, DeepH-E3 stands out as the most stable open-source implementation to date and has been utilized with ABACUS in several example studies DeepH-E (3); Tang et al. (2024); Yang et al. (2024). The ABACUS-DeepH interface is versatile, supporting spin-orbit coupling and magnetic systems, and facilitates ABACUS’s integration with xDeepH–a specialized variant of DeepH for predicting the electronic structures of magnetic materials Li et al. (2023). To address the modified sparsity patterns in hybrid DFT Hamiltonians, a dedicated toolkit, named DeepH-hybrid, has been developed, which verified DeepH’s applicability to hybrid-functional Hamiltonians generated by ABACUS Tang et al. (2024, 2023). An interface between ABACUS and DeepH-DFPT (a generalization of the DeepH approach for deep-learning density functional perturbation theory) is currently under development to accelerate calculations of electron-phonon coupling Li et al. (2024). In addition, a recently developed “Hamiltonian Projection and Reconstruction to atomic Orbitals" (H-PRO) method can transform the DFT Hamiltonian of plane-wave basis into localized basis, making the DeepH compatible with the plane-wave mode of ABACUS Gong et al. (2024). Very recently, a universal materials model (UMM) of DeepH, named DeepH-UMM, has been developed, demonstrating exceptional transferability across a wide range of material structures composed of various elements Wang et al. (2024b). Given ABACUS’s versatility, further development of DeepH-UMM in conjunction with ABACUS holds significant promise for advancing materials discovery.
@@
The current ABACUS-DeepH interface is available on GitHub DeepH . For inference with trained DeepH models, ABACUS provides an efficient method for generating the overlap matrix, from which the Hamiltonian can be predicted by DeepH. It is important to note that the overlap and Hamiltonian matrices are assumed to share the same sparsity pattern in DeepH. To ensure clarity, fixing such sparsity pattern with tools provided in DeepH-hybrid is recommended Tang et al. (2023).
@@ VI.5 DeePTB
DeePTB is an open-source package that leverages deep learning to accelerate ab initio electronic simulations Gu et al. (2024); Zhouyin et al. (2024). Its integration with ABACUS creates a powerful synergy, combining first-principles calculations with advanced machine learning techniques for efficient large-scale electronic structure predictions. Notably, it has demonstrated the capability to simulate systems containing up to millions of atoms Gu et al. (2024), a scale previously unattainable with traditional methods.
@@
This integrated approach significantly enhances the efficiency and scalability of electronic structure simulations, enabling the study of complex and large-scale material systems previously limited by computational constraints. For instance, recent work combining DeePTB with non-equilibrium Green’s function (NEGF) methods has demonstrated unprecedented efficiency in quantum transport simulations of large-scale nanodevicesZou et al. (2024). The ABACUS-DeePTB collaboration demonstrates the potential of combining traditional DFT methods with cutting-edge machine learning in computational materials science, opening new avenues for high-throughput materials discovery and design.
@@
VI.6 PYATB

### v4

@@ V.2 UniPero Model
Wu et al. Wu et al. (2023a) utilized “modular development of deep potential” (ModDP) Wu et al. (2023b) to create a universal interatomic potential for perovskite oxides (UniPero) using a deep neural network with a self-attention mechanism. This potential spans 26 types perovskite oxides involving 200 components and 14 metal elements. All DFT calculations were performed with ABACUS. As depicted in Fig. 28(a), they first used DP-GEN to obtain a converged DPA-1 model Zhang et al. (2024d) for three-element systems such as PbTiO3, SrTiO3. Then, the converged dataset served as a starting point for DP-GEN to improve DPA-1 for four-element perovskite systems such as PbxSr1-xTiO3 and PbZrxTi1-xO3. Ultimately, the final DPA-1 model (UniPero) can describe six-element perovskite systems, including the ternary solid solution Pb(In1/2Nb1/2)O3-Pb(Mg1/3Nb2/3)O3–PbTiO3 (PIN-PMN-PT). UniPero functions as a universal interatomic potential, effectively modeling a wide range of perovskites through molecular dynamics simulations.
@@ V.3 DPA-Semi Model
Liu et al. Liu et al. (2024b) generated over 200,000 DFT data for 19 bulk semiconductors ranging from group IIB to VIA, namely, Si, Ge, SiC, BAs, BN, AlN, AlP, AlAs, InP, InAs, InSb, GaN, GaP, GaAs, CdTe, InTe, CdSe, ZnS, CdS. All of the DFT calculations were performed with the ABACUS v3.2 package with the Perdew-Burke-Ernzerhof (PBE) Perdew, Burke, and Ernzerhof (1996) exchange-correlation functional. The triple-zeta plus double polarization (TZDP) NAO basis sets were used for all DFT calculations. The atomic datasets are adopted as training data to generate an attention-based deep potential model using the DPA-1 method, Zhang et al. (2024d) namely the DPA-Semi model. Fig. 29 shows the procedures for developing the DPA-Semi model.
@@ V.4 DPA-1, DPA-2 and DPA-3
The DPA-1 model Zhang et al. (2024d) is a universal machine learning interatomic potential that leverages the attention mechanism for molecular and materials modeling. Trained on the OC20 benchmark dataset and specialized materials systems, including aluminum-magnesium-copper alloys, high-entropy alloys, and solid-state electrolytes, this architecture achieves robust transferability across diverse atomic environments. With coverage spanning 56 elements in the periodic table, DPA-1 employs a pretraining-fine-tuning paradigm. The attention-based design supports first-principles accuracy while maintaining computational efficiency comparable to classical force fields.
@@
The DPA-2 model Zhang et al. (2024e) employs a multi-task learning framework that effectively integrates datasets from diverse DFT sources. This architecture substantially reduces the model’s dependence on data uniformity, enabling utilization of datasets generated by different methods. Notably, the pretraining phase of DPA-2 incorporates systematic learning of chemical compositions and spatial configurations, which dramatically reduces the demand for task-specific fine-tuning data in downstream applications.
@@
The recently developed DPA-3 Zhang et al. (2025a) constructs a hierarchical graph neural network architecture based on line graph transformations, demonstrating substantial improvements in scaling law adherence and cross-dataset generalization capabilities. During the development of both DPA-2 and DPA-3 iterations, the ABACUS software served as a critical computational infrastructure, facilitating DFT calculations across diverse material datasets. Here are two examples.
@@
First, based on DFT calculations performed using ABACUS, we are developing general machine-learning-based potential models for alloys, covering 53 elements in the periodic table. The elements are Li, Na, K, Be, Mg, Ca, Sr, Sc, Y, La, Ti, Zr, Hf, V, Nb, Ta, Cr, Mo, W, Cu, Ag, Au, Zn, Cd, Mn, Re, Fe, Co, Ni, Os, Ir, Pt, Rh, Ru, Pd, Ce, Pr, Nd, Sm, Gd, Tb, Dy, Ho, Er, Tm, Lu, Al, Ga, In, Si, Ge, Sn, Pb, where bold ones are commonly used in structural alloys. The specific pseudopotentials utilized for these elements are listed in Ref. lin, 2024 and the calculations were performed efficiently by the ABACUS package on the Deep Computing Unit (DCU) hardware. Specifically, the DP-GEN and APEX workflows generated various structures, Zhang et al. (2020c); Li et al. (2025) including perfect and perturbed crystal structures, vacancies, interstitials, and surfaces of metals and alloys. Some structures were selected through the concurrent learning workflow in DP-GEN and then labeled by ABACUS, resulting in a training dataset of $\sim$24,000 entries. By integrating previous training datasets from AI Square and the OpenLAM project, ope (2024a) the latest general LAM for alloys achieved a root mean square error (RMSE) of $\sim$26 meV/atom for energies and $\sim$0.20 eV/Å for atomic forces across the 53 metals and their alloys. This general LAM for alloys demonstrates superior performance in predicting various properties, including lattice parameters, elastic constants, point defects, and surface formation energies, compared to other LAMs. Batatia et al. (2024) ABACUS has proven to be a stable and reliable DFT workhorse for diverse structures covering these 53 elements.
@@
Second, during the construction of the dataset for the DPA-1 and DPA-2 models of high-pressure SuperHydrides, all labels were calculated using ABACUS with plane-wave basis sets. The thermodynamic stability of the discovered structures was confirmed by their energies above the convex hull, which was determined through structure optimization results from ABACUS. Dynamic stability was assessed using the phonon dispersion spectrum calculated by phonopy, Togo et al. (2023); Togo (2023) with all forces calculated by ABACUS. Our DFT calculations were performed on DCU hardware, significantly enhancing the efficiency of SCF calculations compared to traditional CPUs. We obtained a dataset comprising 218,349 data frames, covering 29 elements and a pressure range of 150-250 GPa. Utilizing this data, the DPA-1 model achieved a training RMSE of 48.1 meV/atom for energy and 334.8 meV/Å for force. The DPA-2 model achieved a training RMSE of 55.2 meV/atom for energy and 298.8 meV/Å for force. On the testing dataset, the DPA-1 model yielded an RMSE of 37.6 meV/atom for energy and 171.4 meV/Å for force, while the DPA-2 model yielded an RMSE of 37.4 meV/atom for energy and 122.1 meV/Å for force. Given the large dataset required for the Superhydrides model, ABACUS proved to be the most economical DFT software available. The high-quality training results further attest to the stable performance of ABACUS.
@@ VI.1 DeePKS-kit
DeePKS-kit Chen et al. (2021, 2022a) is an innovative computational framework designed to bridge machine learning with quantum-mechanical simulations. The software integrates deep learning techniques with the Kohn-Sham equations to refine exchange-correlation functionals and can be used with ABACUS or PySCF. Sun et al. (2020) By training neural network on high-fidelity quantum chemistry data, DeePKS-kit dynamically corrects errors in conventional DFT approximations, enabling predictions of electronic properties with near-chemical-accuracy at a fraction of the computational cost.
By adopting both DeePKS-kit and ABACUS packages, we optimize the DeePKS model in two iterative steps. As shown in Fig. 31, $L(\omega)$ is the loss function. Loss function $L(\omega)$ can be composed of various energetic terms including the energy $E$, the atomic force $\mathbf{F}$, the stress tensor $\mathbf{\sigma}$ and the band gap $\varepsilon_{g}$ of the interested system. The first step is the SCF step, where we fix parameters $\{\omega^{\ast}\}$ in the DeePKS model and solve the Kohn-Sham equation to obtain the ground state electronic wave function. The second step is the TRAIN step, where the electronic wave functions are fixed and the model parameters $\{\omega\}$are optimized. These two steps repeat until convergence is reached. Chen et al. (2021)
@@ VI.2 DeePMD-kit
DeePMD Zhang et al. (2018a, b) is a widely-used Wen et al. (2022) machine-learning molecular dynamics method based on the neural-network potential. ABACUS has an interface with the DeePMD-kit package, Wang et al. (2018); Zeng et al. (2023) allowing its deployment as a molecular dynamics engine in executing DeePMD simulations. Trained on system energy, atomic forces, and lattice stress derived from DFT calculations, this neural network potential enables DeePMD to replicate the potential energy surface with first-principles accuracy. Since the neural network potential is computationally more efficient than the DFT method, DeePMD can simulate large systems at a long time scale, which is barely accessible for AIMD simulations.
@@ VI.3 DP-GEN
The DP-GEN framework has proven particularly effective for complex systems like multicomponent alloys, Wu et al. (2021); Shi et al. (2024) chemical reaction pathways, Zeng et al. (2021); Yang et al. (2022) and phase transitions under extreme conditions, Gartner et al. (2020); Fu et al. (2023) where traditional sampling methods struggle to capture rare events or subtle energy landscapes. Open-source and modular by design, DP-GEN supports customization for specific materials or molecules, offering a versatile tool for accelerating the development of robust, quantum-mechanically informed potentials.
@@ VI.4 DeepH
The deep-learning density functional theory Hamiltonian (DeepH) method is a neural network approach based on equivariant graph neural networks for modeling the DFT Hamiltonian as a function of material structure. Li et al. (2022b) Leveraging the sparsity of the Hamiltonian matrix under NAOs and its compatibility with Walter Kohn’s “quantum nearsightness principle”, Kohn (1996) the DeepH method can learn from training data of small structures to infer electronic Hamiltonians of large structures, achieving high prediction accuracy with linear-scaling computational cost. The predicted Hamiltonian may be subsequently utilized for post-processing to evaluate properties including band structures, optical properties, and response properties from density functional perturbation theory, etc. Li et al. (2022b); Gong et al. (2023); Li et al. (2024)
@@
The interface between ABACUS and DeepH was developed in 2022, shortly after the invention of the DeepH approach in 2021. The interface ensures compatibility between ABACUS and various versions of DeepH, including DeepH-pack, Li et al. (2022b); DeepH DeepH-E3, Gong et al. (2023); DeepH-E (3) and DeepH-2. Wang et al. (2024a) Among these, DeepH-E3 stands out as the most stable open-source implementation to date and has been utilized with ABACUS in several example studies. DeepH-E (3); Tang et al. (2024); Yang et al. (2024) The ABACUS-DeepH interface is versatile, supporting spin-orbit coupling and magnetic systems, and facilitates ABACUS’s integration with xDeepH–a specialized variant of DeepH for predicting the electronic structures of magnetic materials. Li et al. (2023) To address the modified sparsity patterns in hybrid DFT Hamiltonians, a dedicated toolkit, named DeepH-hybrid, has been developed, which verified DeepH’s applicability to hybrid-functional Hamiltonians generated by ABACUS. Tang et al. (2024, 2023) An interface between ABACUS and DeepH-DFPT (a generalization of the DeepH approach for deep-learning density functional perturbation theory) is currently under development to accelerate calculations of electron-phonon coupling. Li et al. (2024) In addition, a recently developed “Hamiltonian Projection and Reconstruction to atomic Orbitals" (H-PRO) method can transform the DFT Hamiltonian of plane-wave basis into localized basis, making the DeepH compatible with the plane-wave mode of ABACUS. Gong et al. (2024) Very recently, a universal materials model (UMM) of DeepH, named DeepH-UMM, has been developed, demonstrating exceptional transferability across a wide range of material structures composed of various elements. Wang et al. (2024b) Given ABACUS’s versatility, further development of DeepH-UMM in conjunction with ABACUS holds significant promise for advancing materials discovery.
@@
The current ABACUS-DeepH interface is available on GitHub. DeepH For inference with trained DeepH models, ABACUS provides an efficient method for generating the overlap matrix, from which the Hamiltonian can be predicted by DeepH. It is important to note that the overlap and Hamiltonian matrices are assumed to share the same sparsity pattern in DeepH. To ensure clarity, fixing such sparsity pattern with tools provided in DeepH-hybrid is recommended. Tang et al. (2023)
@@ VI.5 DeePTB
DeePTB is an open-source package that leverages deep learning to accelerate ab initio electronic simulations. Gu et al. (2024); Zhouyin et al. (2024) Its integration with ABACUS creates a powerful synergy, combining first-principles calculations with advanced machine learning techniques for efficient large-scale electronic structure predictions. Notably, it has demonstrated the capability to simulate systems containing up to millions of atoms, Gu et al. (2024) a scale previously unattainable with traditional methods.
@@
This integrated approach significantly enhances the efficiency and scalability of electronic structure simulations, enabling the study of complex and large-scale material systems previously limited by computational constraints. For instance, recent work combining DeePTB with non-equilibrium Green’s function (NEGF) methods has demonstrated unprecedented efficiency in quantum transport simulations of large-scale nanodevices. Zou et al. (2024) The ABACUS-DeePTB collaboration demonstrates the potential of combining traditional DFT methods with cutting-edge machine learning in computational materials science, opening new avenues for high-throughput materials discovery and design.
@@
VI.6 HamGNN
HamGNN (Hamiltonian Graph Neural Network) is an end-to-end model specifically designed for atomic orbital (AO)-based electronic Hamiltonian parameterization. Zhong et al. (2023a) It inherently satisfies the symmetry requirements of rotation, parity, and translation. HamGNN parameterizes spin-orbit coupling (SOC) and magnetic interactions by incorporating physical principles, thereby accurately capturing the SOC effect and effectively describing the magnetic Hamiltonian across various materials. Zhong et al. (2023b) Moreover, HamGNN enhances the accuracy and stability of band structure predictions through a two-stage training strategy: real-space pre-training followed by reciprocal-space fine-tuning, which effectively combines local atomic environments with global electronic behavior within the Brillouin zone. In practical applications—including the modeling of high-order Bi topological insulators, Zhao et al. (2025) Bi ferroelectric domain topological interface states, Luo et al. (2025) and metal-oxide-semiconductor field-effect transistor (MOSFET) device design Wang et al. —HamGNN has demonstrated both high computational efficiency and strong transferability.
The performance of HamGNN relies critically on high-quality training data, where ABACUS plays an essential role. ABACUS stands out as a key tool for building high-precision HSE Hamiltonian training databases, as it is one of the few DFT software packages based on numerical atomic orbitals that support HSE hybrid functionals. When combined with the cross-system transferability of the HamGNN model, this capability enables the development of a universal HSE Hamiltonian model covering the entire periodic table. Zhong et al. (2024a) Recently, the Uni-HamGNN framework achieved universal modeling of SOC effects across the periodic table by decomposing the SOC Hamiltonian into spin-independent terms and SOC correction terms and employing a delta-learning strategy to fit them separately. Zhong et al. (2025) By incorporating a limited number of SOC Hamiltonian matrices from ABACUS, a universal SOC model with HSE-level accuracy can be further established using the delta-learning strategy. This approach, when applied in the training of the universal HSE SOC model, can significantly reduce the expensive computational cost of HSE SOC calculations.
These advancements in universal Hamiltonian modeling open up numerous practical applications with significant computational advantages. Using the universal HamGNN models trained on HSE Hamiltonian matrices as a computational engine can substantially accelerate the computation of carrier mobility and superconductivity-related electron-phonon coupling properties at the hybrid functional level. Zhong et al. (2024b) Furthermore, HamGNN can accurately predict Hamiltonians and non-adiabatic coupling vectors (NACV) that continuously change with atomic displacement, enabling dynamics simulations of excited states at HSE precision. Zhang et al. (2025b) Looking forward, the deepening integration of ABACUS and HamGNN holds promise for developing even more powerful, universal electronic Hamiltonian prediction models, driving the materials computation field towards new horizons of enhanced efficiency and intelligence. The interface for ABACUS and HamGNN is available online. aba (2025)
VI.7 PYATB